# Chiara Arina Centre for Cosmology, Particle Physics and Phenomenology (CP3), Université catholique de Louvain, B-1348 Louvain-la-Neuve, Belgium

### v1

@@
   Jackson Burzynski Simon Fraser University, $8$ University Dr W, Burnaby, BC V5A 1S6, Canada
@@
ARC Centre of Excellence for Dark Matter Particle Physics, Australia
@@ Abstract
preprint: CERN-LPCC-2025-001, IRMP-CP3-25-07, TTK-25-07
@@ II.1 Minimal options and their implementation in high-energy physics software
$\varphi_{\scriptscriptstyle u}$ YS3u $\begin{array}[]{l l l}2000002&2000004&2000006\end{array}$ 00 $({\bf 3},{\bf 1},\frac{2}{3})$
$\varphi_{\scriptscriptstyle d}$ YS3d $\begin{array}[]{l l l}2000001&2000003&2000005\end{array}$ 00 $({\bf 3},{\bf 1},-\frac{1}{3})$
@@
$\varphi_{\scriptscriptstyle\ell}$ YS1e $\begin{array}[]{l l l}2000011&2000013&2000015\end{array}$ 00 $({\bf 1},{\bf 1},-1)$
@@ II.2 Flavoured dark matter: Dark Minimal Flavour Violation
Non-minimally flavoured dark matter can also be introduced to couple to the lepton sector [78, 79, 80], as studied in the context of Dirac and complex scalar DM coupling to right-handed charged leptons. In these constructions, a discrete $\mathbb{Z}_{2}$ symmetry is necessary to stabilise dark matter. Additionally, less minimal models [81] featuring two mediators have been studied. Such a setup can provide an explanation to the longstanding deviations between theoretical predictions and experimental measurements relevant to the anomalous magnetic moment of the muon. While the matrices of couplings $\lambda_{1}$ and $\lambda_{2}$ associated with the two mediators are related by DMFV, the DM spectrum becomes instead free. Finally, flavoured DM models beyond DMFV have also been found to offer a promising avenue for achieving successful baryogenesis [82].
@@ II.4 Towards UV completions - frustrated dark matter
We consider a particular fDM realisation where the mediators are $SU(3)_{C}$ sextets and weak singlets, this last condition being imposed for simplicity, and where the DM is a Dirac fermion. In this scenario, only the scalar mediator directly couples to the SM at tree level, specifically to a pair of quarks of possibly the same electric charge. The hypercharge $Y$ of the mediators depends on the structure of these quark-messenger couplings. While a comprehensive examination of all possible low-dimensional couplings is provided in [88], only one of them is renormalisable. This interaction couples the scalar sextet to a $uu$, $ud$, or $dd$ quark pair pursuant to its hypercharge, and its phenomenological consequences have been explored in [89, 90, 91]. For simplicity, we only consider a sextet scalar with $Y=4/3$ so that the corresponding Lagrangian $\mathcal{L}_{XY}$ [^2] We use the notation $\mathcal{L}_{XY}$ for any Lagrangian including DM-mediator interactions, regardless of the model, for the purpose of uniformity. is given by
[…]
@@ III.1.1 Generalities about the signal of quark-philic dark matter
In the following, we label the contributions from the first to processes by $XX$ and $XY$ respectively, while for the last process we must distinguish the different contributing components. The pair production of a mediator and an anti-mediator can originate from QCD diagrams (labelled by $Y\bar{Y}_{\mathrm{QCD}}$, with a squared matrix element proportional to $\alpha_{s}^{2}$), $t$-channel DM exchange diagrams (labelled by $Y\bar{Y}_{t}$ with a squared matrix element proportional to $\lambda^{4}$), and the corresponding interference (labelled by $Y\bar{Y}_{i}$, with a matrix element proportional to $\alpha_{s}\lambda^{2}$). Conversely, the production of two mediators or two anti-mediators can only be induced by $t$-channel dark matter exchanges in models where the DM state is real, with the corresponding matrix elements being proportional to $\lambda^{4}$. The total new physics cross section $\sigma_{\mathrm{BSM}}$ is thus given, with the dependence on the new physics coupling $\lambda$ factorised and as a function of the new physics masses $M_{X}$ and $M_{Y}$, by:
@@
Specifically, the $XX$ component has a strong dependence on the new physics coupling $\lambda$, to the fourth power, but its modelling requires considering an additional hard visible object in the final state to be detectable. On the other hand, the various contributions involving two mediators or anti-mediators in the final state ($YY$, $Y\bar{Y}$ and $\bar{Y}\bar{Y}$) exhibit different dependencies on $\lambda$. Moreover, these are generally phase-space disfavoured due to the large mediator mass and the fact that we deal with pair production. Eq. (26) also shows that, even in feebly-coupled scenarios, the QCD contribution serves as a baseline component independent of the DM mass. Conversely, other (anti-)mediator pair channels become increasingly relevant with higher $\lambda$ values, and some of them can benefit from parton density enhancements. This is specially relevant for scenarios where DM couples to up or down quarks. In such cases, processes like $uu\to YY$ and $dd\to YY$ can involve a pair of initial valence quarks. Finally, $XY$ production has a weaker $\lambda$ dependence, only to the second power, but is phase-space favoured thanks to the production of a lighter DM state. Consequently, there is no clear model-independent hierarchy between the different contributions, necessitating their inclusion in signal simulations.
In this work, hard-scattering simulations are performed with MadGraph5_aMC@NLO [50] (version 2.9.18) using the MadSTR plugin [43] [^5] The MadSTR plugin can be downloaded from https://code.launchpad.net/~maddevelopers/mg5amcnlo/MadSTRPlugin. to handle resonant contributions appearing at NLO, where an intermediate mediator is produced on-shell and subsequently decays into a DM state and a SM quark. This approach ensures that such contributions are not double-counted across the three new physics processes $XX$, $XY$, and $YY$. For instance, at NLO, the process $pp\to XY\to X(Xj)$ includes configurations where the intermediate mediator is on-shell (to be included in $XY$ simulations) as well as off-shell contributions (to be included in the real-emission corrections to $XX$ production). Furthermore, LO and NLO matrix elements are determined (with MadGraph5_aMC@NLO) using the tree-level Feynman rules, $R_{2}$ contributions, and counterterms relevant to the models introduced in section II and encoded in the UFO format [46, 47] as detailed above. These matrix elements are convolved with the LO and NLO sets of NNPDF4.0 parton densities [126, 127], respectively. Additionally, decays of heavy unstable particles are handled with MadSpin [128] and MadWidth [129] to retain off-shell propagation and spin correlation effects. Finally, parton showering (PS) and hadronisation effects are simulated with Pythia (version 8.306) [130] and matched with fixed-order calculations according to the MC@NLO procedure[131].
@@
define xx = <X state>
define yy = <Y state>
@@
The event generation commands for the $XX$, $XY$, $Y\bar{Y}_{t}$, $YY_{t}$, and $\bar{Y}\bar{Y}_{t}$ processes are as follows:
generate p p > xx xx / <excluded states> [QCD]
generate p p > xx yyy / <excluded states> [QCD]
@@
generate p p > yy yy DMT=2 QCD=0 QED=0 / <excluded states> [QCD]
generate p p > $\sim$ $\sim$ DMT=2 QCD=0 QED=0 / <excluded states> [QCD]
@@
$|\eta(j)|$ $<2.8$ $<2.0$ $<2.8$ $<2.4$ $<2.4$
$p_{T}(j_{1})$ $>150$ GeV $>50$ GeV $>200$ GeV $>30$ GeV $>100$ GeV
$p_{T}(j_{2},...,j_{N_{j}})$ $>30$ GeV $>50$ GeV $>20$ GeV
$E_{T}^{\mathrm{miss}}$ $>200$ GeV -- $>300$ GeV -- $>250$ GeV
$H_{T}^{\mathrm{miss}}$- -- -- $>300$ GeV --
$E_{T}^{\mathrm{miss}}/\sqrt{H_{T}}$- $>5~{}\sqrt{\mathrm{GeV}}$ $>10~{}\sqrt{\mathrm{GeV}}$- --
$\Delta\Phi(j_{i},\vec{p}_{T}^{\mathrm{miss}})$ $>0.4$- $>0.2$ $>0.3$ $>0.5$
$m_{\text{eff}}$- -- $>800$ GeV -- --
$H_{T}$- -- -- $>300$ GeV --
Table 4: Summary of the typical event selection cuts included in the five jets+$E_{T}^{\mathrm{miss}}$ analyses considered. Detailed information can be found in the experimental publications [133, 134, 135, 136, 137].
@@ III.1.2 First generation simplified models
Figure 1: Exclusion limits at 95% confidence level from the reinterpretation of several Run 2 ATLAS and CMS searches [133, 134, 135, 136, 137]. The results are shown for the S3M_uR (left) and S3M_dR (right) real dark matter scenarios described in section II.1, considering two configurations: $\lambda=3.5$ (top row) and $\Gamma_{Y}/M_{Y}=0.05$ (bottom row). For scenarios with $\lambda=3.5$, dotted grey lines represent isolines of constant $\Gamma_{Y}/M_{Y}$ value. Conversely, for scenarios with $\Gamma_{Y}/M_{Y}=0.05$, these lines correspond to isolines of fixed $\lambda$ value. Individual contributions to the bounds are displayed for processes $XX$ (red), $XY$ (green), and $YY$ (dark blue), with the $YY$ process further decomposed into its purely QCD part ($YY_{\rm QCD}$, teal) and its $t$-channel part ($YY_{t}$, turquoise). The yellow gradient highlights regimes where either the perturbative approach becomes increasingly invalid due to large coupling values or the narrow-width approximation loses validity due to a large mediator width-to-mass ratio.
@@
In all the figures, mass configurations to the left of the solid (dashed) orange exclusion line are ruled out at the 95% confidence level, based on state-of-the-art simulations at NLO+PS (LO+PS). For scenarios with $\lambda=3.5$ (top row), the mediator mass is constrained to be larger than approximately $3.2-3.5$ TeV for up-like setups and $2.5-2.7$ TeV for down-like setups, provided that the spectrum compression, defined as $r=M_{Y}/M_{X}-1$, is greater than $0.6-0.7$. This corresponds to an LHC sensitivity to DM masses $M_{X}$ smaller than approximately $2$ TeV and $1.5$ TeV for the up-like and down-like cases, respectively. The loss in sensitivity for decreasing $r$ values (i.e. scenarios closer to the diagonal in the figures) has a twofold origin. First, the $XX$ contribution (red contours), despite being enhanced by the large $\lambda^{4}$ factor as shown in (26), is only significant for very light new physics spectra. Consequently, for heavier mass spectra, such as those close to the current bounds, mediator production (both pair and associated) and subsequent decay drives the constraints. The final-state jets originating from mediator decays thus become softer for increasingly compressed spectra, leading to a reduction in signal selection efficiencies and consequently weaker bounds. Specifically, our predictions show that the LHC currently has no sensitivity to models where the DM mass $M_{X}\gtrsim 2.2$ TeV for up-like setups, and $M_{X}\gtrsim 1.5$ TeV for down-like setups.
@@
In the bottom row of the figure, we address this last issue by considering a second class of scenarios. Here, the two new physics masses $M_{X}$ and $M_{Y}$ remain free parameters, but the coupling $\lambda$ is dynamically computed to enforce a fixed mediator width-to-mass ratio of $\Gamma_{Y}/M_{Y}=0.05$. Consequently, this approach eliminates any issues related to the use of MadSpin for modelling mediator decays. The resulting exclusion contours exhibit a distinct shape compared to the case where $\lambda$ was fixed to $3.5$. For scenarios where the mediator width-to-mass ratio is fixed, the exclusion bounds (orange line) are nearly independent of the dark matter mass, reaching $2.5-3$ TeV and $1.8-2.1$ TeV in the S3M_uR and S3M_dR cases, respectively. As before, the exclusion is primarily driven by the $YY$ channel, except for scenarios involving lighter dark matter masses and a heavy mediator, where associated $XY$ production begins to contribute. This behaviour is consistent with the dependence of the mediator pair production rates on the DM mass, as detailed in the analysis of related matrix elements and cross sections in [25]. Interestingly, the bounds become stronger for increasingly compressed spectra, but this trend arises from the progressively larger values of $\lambda$ required to maintain $\Gamma_{Y}/M_{Y}=0.05$. In such regions of the parameter space, the perturbativity assumption underlying the entire calculation no longer holds so that the exclusion bounds should be interpreted with caution.
@@
Another notable feature of our results is the impact of higher-order QCD corrections on the exclusion bounds. As outlined in [40], NLO QCD corrections influence not only the overall rate of the new physics signal (i.e. the combined contributions from the $XX$, $XY$, and $YY$ channels) but also significantly alter the shapes of key observables, such as the missing transverse momentum $\vec{p}_{T}^{\mathrm{miss}}$ and the missing energy $E_{T}^{\mathrm{miss}}$ used in all LHC analyses considered (see also section III.1.5). As a result, the exclusion contours at LO and NLO are not merely related by a simple translation of each other, but exhibit more complex differences. This is evident in figure 1 for all scenarios considered, where the LO bounds (dashed lines) and NLO bounds (solid lines) can be compared. For S3M_uR models, global NLO effects are generally mild, except for scenarios involving a heavy mediator ($M_{Y}\in[2.5,3.5]$ TeV) and a light dark matter particle ($M_{X}\lesssim 1$ TeV). In this region of parameter space and for decreasing $M_{X}$ values, the $XY$ channel begins to contribute significantly, and it turns out that it exhibits a strong sensitivity to NLO corrections due to the strong coupling already entering in the LO matrix elements. Moreover, while NLO contributions are highly relevant for the $YY_{\mathrm{QCD}}$ and $XX$ channels, these channels contribute negligibly given the mass ranges probed by current LHC analyses. Conversely, QCD corrections to the dominant $YY_{t}$ channel are mild, except for setups with light DM masses, where the dependence of the matrix elements on the DM mass becomes non-trivial [25]. Consequently, NLO contributions influence only a specific portion of the parameter space for S3M_uR models. The situation is markedly different for S3M_dR models. Here, the distinct parton distribution functions (PDFs) involved in the $YY_{t}$ and $XY$ matrix elements lead to much stronger NLO effects for these channels, resulting in exclusion bounds that are significantly more sensitive to higher-order corrections than in the S3M_uR case. This is clearly illustrated in the figures, where shifts in the mediator mass limits approach $500$ GeV for a fixed DM mass below approximately $1$ TeV.
@@
As previously outlined, significant differences are observed between the bounds applicable on S3M_uR and S3M_dR models, despite the overall similarity in the shapes of their exclusion contours. The primary distinction lies in the mass range covered. In the region of parameter space where the current bounds reside, the signal is in both cases predominantly driven by contributions from the $YY_{t}$ and $XY$ channels. These two channels are directly influenced by the flavour of the initial-state quark, with the subprocesses $uu\to YY$ and $ug\to XY$ dominating in the S3M_uR case, and $dd\to YY$ and $dg\to XY$ dominating in the S3M_dR case. These subprocesses are strongly affected by partonic luminosities and potential enhancements from valence quarks. Since the up-quark content in the proton is substantially higher than the down-quark content, the bounds on S3M_dR models are consequently weaker by approximately $500-700$ GeV for a given dark matter mass.
@@
In figure 3, we turn our attention to the F3S class of models, in which the dark matter particle is a real scalar state and the mediator is a vector-like fermion lying in the fundamental representation of $SU(3)_{C}$. We remind that such a setup is typical of composite constructions in theories beyond the Standard Model. Similar to our study of the S3M models above, the dark matter and mediator masses are treated as free parameters, while two complementary approaches are used to define the coupling $\lambda$ of dark matter to quarks. In the first approach, $\lambda$ is fixed to a value of $4.8$, motivated by cosmological considerations (see again section IV). In the second approach, the mediator width-to-mass ratio is set to $\Gamma_{Y}/M_{Y}=0.05$, and the corresponding $\lambda$ value is then dynamically determined. The obtained exclusion contours exhibit a similar shape to those of the S3M models, although they are shifted toward significantly higher mediator mass values. This shift originates from the larger cross sections associated with vector-like quark pair production compared to coloured scalar pair production for a given mediator mass and coupling value [170, 43]. Additionally, scenarios with a fixed coupling value naturally result in higher exclusion limits than in the S3M case due to the increased chosen value for $\lambda$. Consequently, mediator masses as high as $3.5-4$ TeV are excluded for F3S_uR models, and $2.7-3.5$ TeV for F3S_dR models. These bounds are almost insensitive to the dark matter mass, as the relevant matrix elements only depend on $M_{X}$ for much larger dark matter masses. As a result, the $qq\to YY$ subprocess (with $q$ representing either an up- or down-type quark) plays an even more dominant role compared to the S3M models when consider the parameter space as a whole. Enhanced by the high partonic luminosity associated with valence quarks, this subprocess alone is indeed sufficient to drive the exclusion bounds across the entire parameter space, with the only mild exception occurs for dark matter masses $M_{X}\lesssim 500$ GeV.

### v2

@@
   Jackson Burzynski Simon Fraser University, $88$ University Dr W, Burnaby, BC V5A 1S6, Canada
@@ Abstract
preprint: CERN-LPCC-2025-001, IRMP-CP3-25-07, MITP-25-012, TTK-25-07
@@ II.1 Minimal options and their implementation in high-energy physics software
$\varphi_{\scriptscriptstyle u}$ YS3u $\begin{array}[]{l l l}2000002&2000004&2000006\end{array}$ 0 $({\bf 3},{\bf 1},\frac{2}{3})$
$\varphi_{\scriptscriptstyle d}$ YS3d $\begin{array}[]{l l l}2000001&2000003&2000005\end{array}$ 0 $({\bf 3},{\bf 1},-\frac{1}{3})$
@@
$\varphi_{\scriptscriptstyle\ell}$ YS1e $\begin{array}[]{l l l}2000011&2000013&2000015\end{array}$ 0 $({\bf 1},{\bf 1},-1)$
@@ II.2 Flavoured dark matter: Dark Minimal Flavour Violation
Non-minimally flavoured dark matter can also be introduced to couple to the lepton sector [78, 79, 80], as studied in the context of Dirac and complex scalar DM coupling to right-handed charged leptons. In these constructions, a discrete $\mathbb{Z}_{2}$ symmetry is always necessary to stabilise dark matter. Additionally, less minimal models [81] featuring two mediators have been studied. Such a setup can provide an explanation to the longstanding deviations between theoretical predictions and experimental measurements relevant to the anomalous magnetic moment of the muon. While the matrices of couplings $\lambda_{1}$ and $\lambda_{2}$ associated with the two mediators are related by DMFV, the DM spectrum becomes instead free. Finally, flavoured DM models beyond DMFV have also been found to offer a promising avenue for achieving successful baryogenesis [82].
@@ II.4 Towards UV completions - frustrated dark matter
We consider a particular fDM realisation where the mediators are $SU(3)_{C}$ sextets and weak singlets, this last condition being imposed for simplicity, and where the DM is a Dirac fermion. In this scenario, only the scalar mediator directly couples to the SM at tree level, specifically to a pair of quarks of possibly the same electric charge. The hypercharge $Y$ of the mediators depends on the structure of these quark-messenger couplings. While a comprehensive examination of all possible low-dimensional couplings is provided in [88], only one of them is renormalisable. This interaction couples the scalar sextet to a $uu$, udud, or $dd$ quark pair pursuant to its hypercharge, and its phenomenological consequences have been explored in [89, 90, 91]. For simplicity, we only consider a sextet scalar with $Y=4/3$ so that the corresponding Lagrangian $\mathcal{L}_{XY}$ [^2] We use the notation $\mathcal{L}_{XY}$ for any Lagrangian including DM-mediator interactions, regardless of the model, for the purpose of uniformity. is given by
$\mathcal{L}_{XY}=\lambda_{X}\varphi^{\dagger}\bar{X}\psi+\lambda_{qq}\bar{K}_{6}\varphi^{\dagger}\bar{u}^{c}_{R}u_{R}+\text{H.c.}\,$, (11)
@@ III.1.1 Generalities about the signal of quark-philic dark matter
In the following, we label the contributions from the first two processes by $XX$ and $XY$ respectively, while for the last process we must distinguish the different contributing components. The pair production of a mediator and an anti-mediator can originate from QCD diagrams (labelled by $Y\bar{Y}_{\mathrm{QCD}}$, with a squared matrix element proportional to $\alpha_{s}^{2}$), $t$-channel DM exchange diagrams (labelled by $Y\bar{Y}_{t}$ with a squared matrix element proportional to $\lambda^{4}$), and the corresponding interference (labelled by $Y\bar{Y}_{i}$, with a matrix element proportional to $\alpha_{s}\lambda^{2}$). Conversely, the production of two mediators or two anti-mediators can only be induced by $t$-channel dark matter exchanges in models where the DM state is real, with the corresponding matrix elements being proportional to $\lambda^{4}$. The total new physics cross section $\sigma_{\mathrm{BSM}}$ is thus given, with the dependence on the new physics coupling $\lambda$ factorised and as a function of the new physics masses $M_{X}$ and $M_{Y}$, by:
@@
Specifically, the $XX$ component has a strong dependence on the new physics coupling $\lambda$, to the fourth power, but its modelling requires considering an additional hard visible object in the final state to be detectable. On the other hand, the various contributions involving two mediators or anti-mediators in the final state ($YY$, $Y\bar{Y}$ and $\bar{Y}\bar{Y}$) exhibit different dependencies on $\lambda$. Moreover, these are generally phase-space disfavoured due to the large mediator mass required to satisfy LHC constraints on coloured particles and the fact that we deal with pair production. Eq. (26) also shows that, even in feebly-coupled scenarios, the QCD contribution serves as a baseline component independent of the DM mass. Conversely, other (anti-)mediator pair channels become increasingly relevant with higher $\lambda$ values, and some of them can benefit from parton density enhancements. This is specially relevant for scenarios where DM couples to up or down quarks. In such cases, processes like $uu\to YY$ and $dd\to YY$ can involve a pair of initial valence quarks. Finally, $XY$ production has a weaker $\lambda$ dependence, only to the second power, but is phase-space favoured owing to the production of a lighter DM state. Consequently, there is no clear model-independent hierarchy between the different contributions, necessitating their inclusion in signal simulations.
In this work, hard-scattering simulations are performed with MadGraph5_aMC@NLO [50] (version 2.9.18) using the MadSTR plugin [43] [^5] The MadSTR plugin can be downloaded from https://code.launchpad.net/~maddevelopers/mg5amcnlo/MadSTRPlugin. to handle resonant contributions appearing at NLO, where an intermediate mediator is produced on-shell and subsequently decays into a DM state and a SM quark. This approach ensures that such contributions are not double-counted across the three new physics processes $XX$, XYXY, and $YY$. For instance, at NLO, the process $pp\to XY\to X(Xj)$ includes configurations where the intermediate mediator is on-shell (to be included in $XY$ simulations) as well as off-shell contributions (to be included in the real-emission corrections to $XX$ production). Furthermore, LO and NLO matrix elements are determined (with MadGraph5_aMC@NLO) using the tree-level Feynman rules, $R_{2}$ contributions, and counterterms relevant to the models introduced in section II and encoded in the UFO format [46, 47] as detailed above. These matrix elements are convolved with the LO and NLO sets of NNPDF4.0 parton densities [126, 127], respectively. Additionally, decays of heavy unstable particles are handled with MadSpin [128] and MadWidth [129] to retain off-shell propagation and spin correlation effects. Finally, parton showering (PS) and hadronisation effects are simulated with Pythia (version 8.306) [130] and matched with fixed-order calculations according to the MC@NLO procedure[131].
@@
define $x$ = <X state>
define $y$ = <Y state>
@@
The event generation commands for the $XX$, XYXY, $\bar{Y}_{t}, YY$$Y_{t}$, and $\bar{Y}\bar{Y}_{t}$ processes are as follows:
generate p p > $x$ $x$ / <excluded states> [QCD]
generate p p > $x$ yyy / <excluded states> [QCD]
@@
generate p p > $y$ $y$ DMT=2 QCD=0 QED=0 / <excluded states> [QCD]
generate p p > $\sim yy\sim$ DMT=2 QCD=0 QED=0 / <excluded states> [QCD]
@@
$|\eta(j)|$ <2.8<2.8 <2.0<2.0 <2.8<2.8 <2.4<2.4 <2.4<2.4
$p_{T}(j_{1})$ >150>150 GeV >50>50 GeV >200>200 GeV >30>30 GeV >100>100 GeV
$p_{T}(j_{2},...,j_{N_{j}})$ >30>30 GeV >50>50 GeV >20>20 GeV
$E_{T}^{\mathrm{miss}}$ >200>200 GeV - >300>300 GeV - >250>250 GeV
$H_{T}^{\mathrm{miss}}$ - - - >300>300 GeV -
$E_{T}^{\mathrm{miss}}/\sqrt{H_{T}}$ - $5~\sqrt{\mathrm{GeV}} >$$10~\sqrt{\mathrm{GeV}} - -$
$\Delta\Phi(j_{i},\vec{p}_{T}^{\mathrm{miss}})$ >0.4>0.4 - >0.2>0.2 >0.3>0.3 >0.5>0.5
$m_{\text{eff}}$ - - >800>800 GeV - -
$H_{T}$ - - - >300>300 GeV -
[…]
@@ III.1.2 First generation simplified models
Figure 1: Exclusion limits at 95% confidence level from the reinterpretation of several Run 2 ATLAS and CMS searches [133, 134, 135, 136, 137]. The results are shown for the S3M_uR (left) and S3M_dR (right) real dark matter scenarios described in section II.1, considering two configurations: $\lambda=3.5$ (top row) and $\Gamma_{Y}/M_{Y}=0.05$ (bottom row). For scenarios with $\lambda=3.5$, dotted grey lines represent isolines of constant $\Gamma_{Y}/M_{Y}$ value. Conversely, for scenarios with $\Gamma_{Y}/M_{Y}=0.05$, these lines correspond to isolines of fixed $\lambda$ value. Individual contributions to the bounds are displayed for processes XXXX (red), XYXY (green), and YYYY (dark blue), with the $YY$ process further decomposed into its purely QCD part ($YY_{\rm QCD}$, teal) and its $t$-channel part ($YY_{t}$, turquoise). The yellow gradient highlights regimes where either the perturbative approach becomes increasingly invalid due to large coupling values or the narrow-width approximation loses validity due to a large mediator width-to-mass ratio.
@@
In all the figures, mass configurations to the left of the solid (dashed) orange exclusion line are ruled out at the 95% confidence level, based on state-of-the-art simulations at NLO+PS (LO+PS). For scenarios with $\lambda=3.5$ (top row), the mediator mass is constrained to be larger than approximately 3.23.53.2-3.5 TeV for up-like setups and 2.52.72.5-2.7 TeV for down-like setups, provided that the spectrum compression, defined as $r=M_{Y}/M_{X}-1$, is greater than 0.60.70.6-0.7. This corresponds to an LHC sensitivity to DM masses $M_{X}$ smaller than approximately $2$ TeV and $1.5$ TeV for the up-like and down-like cases, respectively. The loss in sensitivity for decreasing $r$ values (i.e. scenarios closer to the diagonal in the figures) has a twofold origin. First, the $XX$ contribution (red contours), despite being enhanced by the large $\lambda^{4}$ factor as shown in (26), is only significant for very light new physics spectra. Consequently, for heavier mass spectra, such as those close to the current bounds, mediator production (both pair and associated) and subsequent decay drives the constraints. The final-state jets originating from mediator decays thus become softer for increasingly compressed spectra, leading to a reduction in signal selection efficiencies and consequently weaker bounds. Specifically, our predictions show that the LHC currently has no sensitivity to models where the DM mass $M_{X}\gtrsim 2.2$ TeV for up-like setups, and $M_{X}\gtrsim 1.5$ TeV for down-like setups.
@@
In the bottom row of the figure, we address this last issue by considering a second class of scenarios. Here, the two new physics masses $M_{X}$ and $M_{Y}$ remain free parameters, but the coupling $\lambda$ is dynamically computed to enforce a fixed mediator width-to-mass ratio of $\Gamma_{Y}/M_{Y}=0.05$. Consequently, this approach eliminates any issues related to the use of MadSpin for modelling mediator decays. The resulting exclusion contours exhibit a distinct shape compared to the case where $\lambda$ was fixed to $3.5$. For scenarios where the mediator width-to-mass ratio is fixed, the exclusion bounds (orange line) are nearly independent of the dark matter mass, reaching 2.532.5-3 TeV and 1.82.11.8-2.1 TeV in the S3M_uR and S3M_dR cases, respectively. As before, the exclusion is primarily driven by the $YY$ channel, except for scenarios involving lighter dark matter masses and a heavy mediator, where associated $XY$ production begins to contribute. This behaviour is consistent with the dependence of the mediator pair production rates on the DM mass, as detailed in the analysis of related matrix elements and cross sections in [25]. Interestingly, the bounds become stronger for increasingly compressed spectra, but this trend arises from the progressively larger values of $\lambda$ required to maintain $\Gamma_{Y}/M_{Y}=0.05$. In such regions of the parameter space, the perturbativity assumption underlying the entire calculation no longer holds so that the exclusion bounds should be interpreted with caution.
@@
Another notable feature of our results is the impact of higher-order QCD corrections on the exclusion bounds. As outlined in [40], NLO QCD corrections influence not only the overall rate of the new physics signal (i.e. the combined contributions from the $XX$, XYXY, and $YY$ channels) but also significantly alter the shapes of key observables, such as the missing transverse momentum $\vec{p}_{T}^{\mathrm{miss}}$ and the missing energy $E_{T}^{\mathrm{miss}}$ used in all LHC analyses considered (see also section III.1.5). As a result, the exclusion contours at LO and NLO are not merely related by a simple translation of each other, but exhibit more complex differences. This is evident in figure 1 for all scenarios considered, where the LO bounds (dashed lines) and NLO bounds (solid lines) can be compared. For S3M_uR models, global NLO effects are generally mild, except for scenarios involving a heavy mediator ($M_{Y}\in[2.5,3.5]$ TeV) and a light dark matter particle ($M_{X}\lesssim 1$ TeV). In this region of parameter space and for decreasing $M_{X}$ values, the $XY$ channel begins to contribute significantly, and it turns out that it exhibits a strong sensitivity to NLO corrections due to the strong coupling already entering in the LO matrix elements. Moreover, while NLO contributions are highly relevant for the $YY_{\mathrm{QCD}}$ and $XX$ channels, these channels contribute negligibly given the mass ranges probed by current LHC analyses. Conversely, QCD corrections to the dominant $YY_{t}$ channel are mild, except for setups with light DM masses, where the dependence of the matrix elements on the DM mass becomes non-trivial [25]. Consequently, NLO contributions influence only a specific portion of the parameter space for S3M_uR models. The situation is markedly different for S3M_dR models. Here, the distinct parton distribution functions (PDFs) involved in the $YY_{t}$ and $XY$ matrix elements lead to much stronger NLO effects for these channels, resulting in exclusion bounds that are significantly more sensitive to higher-order corrections than in the S3M_uR case. This is clearly illustrated in the figures, where shifts in the mediator mass limits approach $500$ GeV for a fixed DM mass below approximately $1$ TeV.
@@
As previously outlined, significant differences are observed between the bounds applicable on S3M_uR and S3M_dR models, despite the overall similarity in the shapes of their exclusion contours. The primary distinction lies in the mass range covered. In the region of parameter space where the current bounds reside, the signal is in both cases predominantly driven by contributions from the $YY_{t}$ and $XY$ channels. These two channels are directly influenced by the flavour of the initial-state quark, with the subprocesses $uu\to YY$ and $ug\to XY$ dominating in the S3M_uR case, and $dd\to YY$ and $dg\to XY$ dominating in the S3M_dR case. These subprocesses are strongly affected by partonic luminosities and potential enhancements from valence quarks. Since the up-quark content in the proton is substantially higher than the down-quark content, the bounds on S3M_dR models are consequently weaker by approximately 500700500-700 GeV for a given dark matter mass.
@@
In figure 3, we turn our attention to the F3S class of models, in which the dark matter particle is a real scalar state and the mediator is a vector-like fermion lying in the fundamental representation of $SU(3)_{C}$. We remind that such a setup is typical of composite constructions in theories beyond the Standard Model. Similar to our study of the S3M models above, the dark matter and mediator masses are treated as free parameters, while two complementary approaches are used to define the coupling $\lambda$ of dark matter to quarks. In the first approach, $\lambda$ is fixed to a value of $4.8$, motivated by cosmological considerations (see again section IV). In the second approach, the mediator width-to-mass ratio is set to $\Gamma_{Y}/M_{Y}=0.05$, and the corresponding $\lambda$ value is then dynamically determined. The obtained exclusion contours exhibit a similar shape to those of the S3M models, although they are shifted toward significantly higher mediator mass values. This shift originates from the larger cross sections associated with vector-like quark pair production compared to coloured scalar pair production for a given mediator mass and coupling value [170, 43]. Additionally, scenarios with a fixed coupling value naturally result in higher exclusion limits than in the S3M case due to the increased chosen value for $\lambda$. Consequently, mediator masses as high as 3.543.5-4 TeV are excluded for F3S_uR models, and 2.73.52.7-3.5 TeV for F3S_dR models. These bounds are almost insensitive to the dark matter mass, as the relevant matrix elements only depend on $M_{X}$ for much larger dark matter masses. As a result, the $qq\to YY$ subprocess (with $q$ representing either an up- or down-type quark) plays an even more dominant role compared to the S3M models when consider the parameter space as a whole. Enhanced by the high partonic luminosity associated with valence quarks, this subprocess alone is indeed sufficient to drive the exclusion bounds across the entire parameter space, with the only mild exception occurs for dark matter masses $M_{X}\lesssim 500$ GeV.